# What Matters in Learning Facts in Language Models? Multifaceted Knowledge Probing with Diverse Multi-Prompt Datasets

### v1

@@ 5.4 Can BELIEFs mitigate bias?
We measure bias in both single- and multi-prompt evaluations. We use average bias-level over all relational templates to represent prompt bias existing during the single-prompt evaluation. Then, we simulate the prompt sampling for accuracy calculation in multi-prompt evaluation by averaging output distributions over templates of the same relation. We then use the bias-level of this average distribution to quantify bias in multi-prompt evaluation. In experiments with BERT models, we observe consistent improvement with multi-prompt evaluation. For example, BERTwwm has average bias-level of 0.304 for single- and 0.217 for multi-prompt evaluations. Taking P31:instance-of [^13] https://www.wikidata.org/wiki/Property:P31 as an example, the average probability of “science” over all templates is 8.30%, but it rises to 52.79% for template: “[Y] contains [X] as one of its elements.”
6 Differentiating PLMs in Fact Probing
In this section, we compare the performance of different PLMs and explore the factors that impact factual knowledge learning during pre-training.
6.1 Factors affecting the acquisition of facts
1) Pre-training strategy From Table 3, we can observe that BERTwwm outperforms BERTlarge in terms of all metrics, while BERTwwm differs from BERTlarge only in the masking strategy during pre-training. The superiority of BERTwwm likely stems from its challenging pre-training paradigm, which requires recalling whole words without sub-token information, enhancing word-level contextual understanding. This underscores the importance of pre-training strategy in knowledge acquisition.
[…]
range stdev
@@ 6.1 Factors affecting the acquisition of facts
2) Model size: Then, we compare models with difference sizes as shown in Table 5.141414Owing to the high computational cost of inference on large LLMs like Llama2-70B, we select only five manually rewritten templates with 4-template ICL setting for evaluation. We find that larger LLMs consistently achieve higher accuracy in predicting facts. Combining with the improvement from BERTbase to BERTlarge from Table 3, the importance of model size in fact acquisition during pre-training is confirmed. Moreover, larger models also demonstrate better robustness in handling diverse prompts and shows better reliability.
PLMs Acc@1 Fluctuation Consist Overconf 1-word
range stdev gen.
Llama2-7B .3384 .2602 .0299 .1269 -.0731 .4752
Llama3-8B .3427 .2864 .0350 .0240 -.1119 .1572
Phi3-mini .3498 .2374 .0292 .1465 .1758 .8641
Table 6: BELIEF-ICL evaluation with zero-shot setting.
3) Pre-training corpora: Table 5 shows that Llama3-8B outperforms the larger Llama2-13B in fact probing. This is likely due to Llama3’s pre-training corpus being seven times larger than Llama2. Additionally, Llama3-70B surpasses Llama2-70B, confirming the importance of pre-training data volume for fact acquisition.
In the zero-shot evaluation using the entire MyriadLAMA, as shown in Table 6, Phi3-mini outperforms Llama2-7B and Llama3-8B in knowledge retrieval and fidelity to instructions (one-word generation rate). Given that Phi3-mini has about half the model size of Llama2-7B and Llama3-8B (3.8B), and model size typically enhances knowledge retrieval, this result is notable. This superior performance can be attributed to the high-quality, textbook-like material used for pre-training the Phi3 models, highlighting the significant impact of high-quality training data.
@@ 6.2 Factors affecting the reliability
Table 3 shows a significant difference in Overconf between BERT models and Llama3-8B, with BERT models being overconfident and Llama3-8B being underconfident. In this section, we explore the reasons for these differences and investigate additional factors affecting reliability beyond model size.
1) The number of output tokens: One main difference in Overconf calculation between encoder- and decoder-based PLMs is that the decoder-based PLMs will generate multiple tokens. Thus, we investigate the effect of output token count on Overconf values. We divide the MyriadLAMA prompt set into groups based on the number of tokens generated. For each group, we calculate the probability of the entire token sequence and compute Overconf for token counts from 1 to 5.151515The prompts generated within five tokens cover 98.78% of Llama3-8B’s generations with the 4-template ICL setting.
The Overconf values for each group of 1 to 5 output tokens on Llama3-8B (4-template) are -0.1030, -0.0906, -0.0297, -0.0546, and 0.0573, showing models become more overconfident with more output tokens. This trend is consistent across models.
2) The writing style of training data: As shown in Table 5, the Phi3 and BERT models, which were pre-trained on textbook-like text such as Wikipedia, exhibit significant overconfidence in their predictions. In contrast, the Llama2 and Llama3 models, which use general online data, show better calibration between accuracy and confidence (Figure 1 Right). This is probably because the Phi3 and BERT models are trained to predict textbook-like text with more expressions directly describing factual knowledge (e.g., definition sentences). This highlights the significant impact of the pre-training corpora’s writing style on PLM reliability.
@@ 6.3 How do PLMs perceive facts differently?
The results are shown in Figure 2. The average sharing rate among BERT models is 69.07%, while for Llama2-7B, Llama3-8B, and Phi3-mini in the zero-shot ICL setting, it is 67.58%. In comparison, the average sharing rate between encoder- and decoder-based PLMs reduced to 47.10%. Additionally, knowledge sharing rates in both zero-shot and 4-template settings indicate that incorporating context examples improves the knowledge elicited by the models. However, about 10% of knowledge can still only be elicited in the zero-shot setting.
@@ 8 Related work
The LAMA probe was first proposed to evaluate the utility of PLMs as knowledge bases via solving the fill-in-the-blank task Petroni et al. (2019). Several researchers extend the LAMA probe to evaluate PLMs’ ability to understand facts from diverse linguistic aspects, such as the effect of negation/mispriming Kassner and Schütze (2020), distractors Pandia and Ettinger (2021), multilingual understanding Keleg and Magdy (2023); Zhao et al. (2024) and models’ consistency facing prompts with minor nuances Fierro and Søgaard (2022); Elazar et al. (2021). However, these studies lack the inspection of PLMs’ reliability in knowledge prediction, which is vital in deploying LLMs to real-world tasks. Moreover, solving the fill-in-the-blank task by LLMs with the causal LM objective can underestimate their knowledge understanding ability.
@@ 9 Conclusions
This paper presents multi-faceted factual probing frameworks, BELIEF and BELIEF-ICL, for encoder-based PLMs and decoder-based LLMs, respectively. Leveraging a multi-prompt dataset, BELIEFs provides various evaluation metrics, including accuracy, consistency, and reliability, enabling a thorough assessment of PLMs’ comprehension of factual knowledge. To make BELIEF more reliable, we build a new multi-prompt dataset for knowledge probing, MyriadLAMA, featuring diverse prompts for each fact. We conducted extensive experiments of multiple encoder-based PLMs and recent LLMs. Based on the evaluation results, we verify the effectiveness of BELIEF in accurately and comprehensively probing knowledge by revealing differences between PLMs and demonstrating its ability to mitigate bias. Additionally, we identify key factors affecting the correctness and reliability of PLMs’ answers in factual probing, such as model size, pre-training strategy and corpora. Finally, by probing facts in different ways, we find that PLMs hold more knowledge than what is revealed by using the optimal template, highlighting the limitations of prompt-based factual probing.
@@ 10 Limitations
MyriadLAMA contains an extensive amount of prompts, which leads to high evaluation costs. In the future, we aim to extract a diverse yet robust subset from MyriadLAMA to enable a more efficient evaluation of factual knowledge.
@@ A.1 The extension of entities
LAMA-UHN is a refined subset derived from the LAMA dataset, which LAMA originates from T-REx Elsahar et al. (2018). T-REx is a large-scale knowledge base containing 11 million real-world knowledge triples, aligned with 3.09 million Wikipedia abstracts, designed to create large-scale alignments between Wikipedia abstracts and Wikidata triples. To achieve this alignment, T-REx employed three distinct aligners—NoSub, AllEnt, and SPO—each offering varying levels of accuracy (0.98, 0.96, and 0.88, respectively) as measured on a test set. Despite the high alignment accuracy of all three aligners, LAMA-UHN selects only the triples aligned by NoSub, the aligner with the highest accuracy. While this choice ensures the high correctness of triples within LAMA, it potentially compromises the ability to fairly assess a PLM’s capability in understanding facts, as it may overlook valid answers during evaluation. To address this limitation, we expand the MyriadLAMA dataset by incorporating triples aligned by all three aligners—NoSub, AllEnt, and SPO—found in T-REx, based on the subject-relation pairs present in LAMA-UHN. As the result, we increase the number of unique triples from 27,106 to 34,048 as shown in Table 1.
@@ A.3 Example of extended relational templates in MyriadLAMA
We display part of the created templates in MyriadLAMA. We randomly select two manually rewritten templates and three auto-generated templates of these two templates. We show the sampled templates for all the relations in Tables 1920212223 and 24.
@@ Appendix B The Advangage of MyriadLAMA
Given that our study seeks to mitigate the influence of individual prompt bias in evaluations, the availability of a wide range of prompts characterized in both quantity and diversity is crucial. The diversity ensures that different prompts can capture different aspects of the true knowledge distribution. On the other side, the quality or correctness of prompts ensure the evaluation can accurately reflect the true ability of understanding facts.
@@ D.2 Evaluation
ICL settings Fact prediction (Acc@1) One-word generation
@@ F.1 Model cards
LLMs Architecture Model size Pre-training corpora
@@
[…]
BERTlarge Encoder-based 336M 3.3G
BERTwwm Encoder-based 336M 3.3G
ALBERTbase Encoder-based* 11.8M 3.3G
ALBERTlarge Encoder-based* 223M 3.3G
[…]
Llama2-13B Decoder-based 13B 2.0T
Llama2-70B Decoder-based 70B 2.0T
[…]
Llama3-70B Decoder-based 70B 15T+
[…]
Phi3-small Decoder-based 7B 4.8T
Phi3-medium Decoder-based 14B 4.8T
*ALBERT use parameter compression techniques, including parameter sharing between token embeddings and transformer layers.
@@ F.3 Evaluation results on all PLMs based on BELIEFs
We present all evaluation results and their computational costs in this section. In Table 16, we report the full-scale experiments using all the prompts provided by MyriadLAMA. This includes PLMs with 8B parameters or fewer, such as BERTbase, BERTlarge, BERTwwm, ALBERTbase, ALBERTlarge, Llama2-7B, Llama3-8B, Phi3-mini (3.8B), and Phi3-small (7B). For decoder-based models, we conduct experiments on 4 types of ICL settings.
@@
Furthermore, we display the calibration level between accuracy and confidence for a straightforward inspection on Overconf metrics. We show the calibration figures of models with full-scale experiments in Figure 5 and experiments with partial prompts in Figure 6.
@@
[…]
4-relation .6286 .1221 .0150 .3753 -.1335 .9060
4-template .6616 .0294 .0036 .4163 -.0933 .9299
@@
[…]
4-relation .6871 .1236 .0156 .3659 -.0783 .9071
4-template .7268 .0220 .0026 .4015 -.0582 .9187
@@
(d) Llama3-8B with 4 ICL settings
(e) Phi3-mini with 4 ICL settings
(f) Phi3-small with 4 ICL settings
@@
(b) Llama2-13B with 2 ICL settings
(c) Llama2-70B with 2 ICL settings
(d) Llama3-8B with 2 ICL settings
(e) Llama3-70B with 2 ICL settings
(f) Phi3-mini with 4 ICL settings
(g) Phi3-small with 4 ICL settings
(h) Phi3-medium with 2 ICL settings

### v3

@@ 6.1 Factors affecting the recall accuracy
range SD
@@
2) Model size. Table 5 compares the knowledge recall abilities of LLMs with difference sizes. [^14] Owing to the high computational cost of inference on large LLMs like Llama2-70B, we select only five manually rewritten templates with 4-template ICL setting for evaluation. We can observe that larger LLMs consistently achieve higher accuracy in predicting facts. Combining with the improvement from BERTbase to BERTlarge from Table 3, the importance of model size in fact acquisition during pre-training is confirmed.
[…]
range SD
Llama2-7B-IT .2925 .1980 .0253 .1151 .2605
Llama3-8B-IT .3578 .2213 .0262 .1660 .1402
Phi3-mini .4258 .2437 .0292 .1782 .2171
Table 6: BELIEF-ICL evaluation with zero-shot setting on instruction-tuned LLMs.
3) Pre-training corpora. Table 5 shows that Llama3-8B outperforms larger Llama2-13B in fact probing. This is likely due to Llama3’s pre-training corpus being seven times larger than Llama2 (Table 2). Meanwhile, Llama3-70B surpasses Llama2-70B, confirming the importance of pre-training data volume for fact acquisition.
In the zero-shot evaluation using the entire MyriadLAMA, as shown in Table 6, Phi3-mini outperforms Llama2-7B-IT and Llama3-8B in knowledge retrieval. Given that Phi3-mini (3.8B) has about half the model size of Llama2-7B-IT and Llama3-8B-IT, and model size typically enhances knowledge retrieval, this result is notable. This superior performance can be attributed to the high-quality, textbook-like material used for pre-training the Phi3 models, highlighting the significant impact of high-quality training data.
4) Instruction-tuning. Table 7 confirms that instruction-tuned Llama2-7B-IT exhibit a higher one-word generation rate than Llama2-7B, as expected. However, the instruction-tuned LLM consistently demonstrate lower Acc@1 scores on different ICL settings. This indicates a potential negative impact of instruction-tuning on the models, where general language understanding can improve, but some factual knowledge is partially lost as a result of the tuning process.
5) Inclusion and selection of demonstrations. As shown in Table 7 and Table 18, using demonstrations in prompts consistently improves Acc@1. Including few-shot demonstrations with same templates to the target question can nearly double Acc@1 values (from zero-shot to 4-template settings). Closer demonstrations also enhance performance across all metrics, highlighting a significant gap between the factual knowledge LLMs memorized and what they can actually recall.
[…]
range SD
Llama2-7B 0-shot .3385 .2602 .0299 .1269 -.1119 .4752
4-rand. .4816 .2250 .0270 .2312 -.0894 .8247
4-rel. .6286 .1221 .0150 .3753 -.1335 .9060
4-templ. .6616 .0294 .0036 .4163 -.0933 .9299
Llama2-7B-IT 0-shot .2925 .1980 .0253 .1151 .2605 .9069
4-rand. .4334 .1958 .0229 .2128 .2410 .9081
4-rel. .5576 .0791 .0092 .3341 .1900 .9314
4-templ. .5896 .0439 .0050 .3687 .2061 .9380
Table 7: Evaluation results of pretrained Llama2-7B (above) and instructed-tuned Llama2-7B-IT (below).
@@ 6.2 Factors affecting the reliability
Table 3 shows a significant difference in Ovconf between BERT models and Llama3-8B, with BERT models being overconfident and Llama3-8B being underconfident. In this section, we explore the reasons for these differences and investigate additional factors affecting reliability beyond model size.
1) The number of output tokens. One main difference in Ovconf calculation between encoder- and decoder-based PLMs is that the decoder-based PLMs will generate multiple tokens. Thus, we investigate the effect of output token count on Ovconf values. We divide the MyriadLAMA prompt set into groups based on the number of tokens generated. For each group, we calculate the probability of the entire token sequence and compute Ovconf for token counts from 1 to 5.151515The prompts generated within five tokens cover 98.78% of Llama3-8B’s generations with the 4-template ICL setting.
The Ovconf values for each group of 1 to 5 output tokens on Llama3-8B (4-template) are -0.1030, -0.0906, -0.0297, -0.0546, and 0.0573, showing models become more overconfident with more output tokens. This trend is consistent across models.
Figure 3: Calibration plots of Llama2-7B (left) and Llama2-7B-IT (right).)
2) Instruction-tuning inflates LLMs confidence. Table 7 and Figure 3 confirm that instruction-tuned LLMs make the models overly confident in their outputs. The pretraining use more diverse language data with uncertainties, which can lead to a more calibrated output confidence. Instruction-tuning narrow the LLMs’ exposure to specific tasks, reducing its ability to express uncertainty and making it more likely to provide overconfident outputs.
3) Model size: Large models consistently demonstrate improved reliability, as illustrated in Table 5.
6.3 Factors affecting the robustness
1) Larger model cannot make zero-shot knowledge prompt more robust. Similar to accuracy and reliability, few-shot knowledge prompts show improved robustness against accuracy fluctuations and consistency as model size increases. However, this effect is absent in zero-shot settings. For instance, the SD for the Llama2 family are 0.2014, 0.2131, and 0.2126 for the 7B, 13B, and 70B models, respectively. Similar inconsistencies are observed across other LLM families with varying model sizes. Refer Table 18 for more details.
@@ 6.3 Factors affecting the robustness
2) Instruction-tuning making fluctuation less influenced by context. Table 7 and Table 17 show that the instruction-tuned models exhibit reduced fluctuation (smaller range and SD) in the zero-shot, 4-random, and 4-relation ICL settings, but perform worse in the 4-template setting. This suggests that instruction-tuned models become less influenced by context and more reliant on the instruction itself.
In contrast, the Consist measure consistently decreased in instruction-tuned models, suggesting that while instruction-tuning improves instruction interpretation, it may weaken semantic understanding, especially with paraphrases.
@@ 6.4 How do PLMs perceive facts differently?
Figure 4 shows the results. The average sharing rate among BERT models is 69.1%, while it is 68.7% for Llama2-7B, Llama2-7B-IT, and Phi3-mini in the zero-shot setting. In comparison, the average sharing rate between encoder- and decoder-based PLMs reduced to 47.1%. Meanwhile, knowledge sharing rates in both zero-shot and 4-template settings indicate that incorporating examples improves the knowledge elicited by the PLMs. However, about 10% of knowledge can still only be elicited in the zero-shot setting (see red boxes).
@@ 8 Related Work
The LAMA probe was first proposed to evaluate the utility of PLMs as knowledge bases via solving the fill-in-the-blank task Petroni et al. (2019). Several researchers extend the LAMA probe to evaluate PLMs’ ability to understand facts from diverse linguistic aspects, such as the effect of negation/mispriming Kassner and Schütze (2020), distractors Pandia and Ettinger (2021), multilingual understanding Keleg and Magdy (2023); Zhao et al. (2024) and models’ consistency facing prompts with minor nuances Fierro and Søgaard (2022); Elazar et al. (2021). However, these studies lack the inspection of PLMs’ reliability in knowledge prediction, which is vital in deploying LLMs to real-world tasks. Moreover, solving the fill-in-the-blank task by LLMs with the causal LM objective can underestimate their knowledge recall ability.
@@ 9 Conclusions
This paper presents the multi-faceted factual probing benchmarks, BELIEF and BELIEF-ICL, for encoder- and decoder-based PLMs, respectively. Leveraging a multi-prompt dataset, BELIEFs provide various evaluation metrics, including accuracy, consistency, and reliability, enabling a thorough evaluation of PLMs’ knowledge recall abilities. To make BELIEFs more reliable, we build a new multi-prompt dataset for knowledge probing, MyriadLAMA, featuring diverse prompts for each fact. We conducted extensive experiments of multiple encoder-based PLMs and recent LLMs.
Based on the evaluation results, we identify key factors affecting the accuracy, reliability and robustness of PLMs’ fact recall, such as model size, pre-training strategy and corpora, and ICL settings. We also reveal the negative effect of instruction-tuning in recall factual knowledge from LLMs. This highlights the need for careful design of instruction-tuning to preserve LLMs’ knowledge recall abilities. Finally, by probing facts in different ways, we find that PLMs hold more knowledge than what is revealed by using the optimal template, highlighting the limitations of prompt-based factual probing.
@@ 10 Limitations
MyriadLAMA contains an extensive amount of prompts, which leads to high evaluation costs. In the future, we aim to extract a diverse yet robust subset from MyriadLAMA to enable a more efficient evaluation of factual knowledge. Additionally, MyriadLAMA is built upon LAMA-UHN, which includes only 41 relationships. Expanding the range of relations is essential to improve coverage in the evaluation of factual knowledge. Lastly, we need to evaluate closed-source LLMs, such as GPT-4 and Claude, to examine performance differences between them and open-source LLMs.
Acknowledgements
This work was partially supported by the special fund of Institute of Industrial Science, The University of Tokyo, by JSPS KAKENHI Grant Number JP21H03494, and by JST, CREST Grant Number JPMJCR19A4, Japan.
@@ A.1 The extension of entities
LAMA-UHN is a refined subset derived from the LAMA dataset, which LAMA originates from T-REx Elsahar et al. (2018). T-REx is a large-scale knowledge base containing 11 million real-world knowledge triples, aligned with 3.09 million Wikipedia abstracts, designed to create large-scale alignments between Wikipedia abstracts and Wikidata triples. To achieve this alignment, T-REx employed three distinct aligners—NoSub, AllEnt, and SPO—each offering varying levels of accuracy (0.98, 0.96, and 0.88, respectively) as measured on a test set. Despite the high alignment accuracy of all three aligners, LAMA-UHN selects only the triples aligned by NoSub, the aligner with the highest accuracy. While this choice ensures the high correctness of triples within LAMA, it potentially compromises the ability to fairly assess a PLM’s knowledge recall ability, as it may overlook valid answers during evaluation. To address this limitation, we expand the MyriadLAMA dataset by incorporating triples aligned by all three aligners—NoSub, AllEnt, and SPO—found in T-REx, based on the subject-relation pairs present in LAMA-UHN. As the result, we increase the number of unique triples from 27,106 to 34,048 as shown in Table 1.
@@ A.3 Example of extended relational templates in MyriadLAMA
We display part of the created templates in MyriadLAMA. We randomly select two manually rewritten templates and three auto-generated templates of these two templates. We show the sampled templates for all the relations in Tables 2021222324 and 25.
@@ Appendix B The Advangage of MyriadLAMA
Given that our study seeks to mitigate the influence of individual prompt bias in evaluations, the availability of a wide range of prompts characterized in both quantity and diversity is crucial. The diversity ensures that different prompts can capture different aspects of the true knowledge distribution. On the other side, the quality or correctness of prompts ensure the evaluation can accurately reflect the true knowledge recall ability.
@@ D.2 Evaluation
ICL settings Fact prediction (Acc@1) 1-word ratio
@@ F.1 Model cards
LLMs Architecture IT† Model size Pre-training corpora
@@
BERTbase Encoder-based No 110M $\left.\begin{array}[]{@{}r@{}}\text{3.3B words}\\ \text{3.3B words}\\ \text{3.3B words}\\ \text{3.3B words}\\ \text{3.3B words}\end{array}\right\}$ BookCorpus (11,038 unpublished books) and English Wikipedia (excluding lists, tables, and headers)
BERTlarge Encoder-based No 336M
BERTwwm Encoder-based No 336M
ALBERTbase Encoder-based* No 11.8M
ALBERTlarge Encoder-based* No 223M
Llama2-7B Decoder-based No 7B $\left.\begin{array}[]{@{}r@{}}\text{2.0T tokens}\\ \text{2.0T tokens}\\ \text{2.0T tokens}\\ \text{2.0T tokens}\\ \text{2.0T tokens}\\ \text{2.0T tokens}\end{array}\right\}$ Publicly available online data (excluding sites containing personal info; factual knowledge sources are upsampled)
Llama2-13B Decoder-based No 13B
Llama2-70B Decoder-based No 70B
Llama2-7B-IT Decoder-based Yes 13B
Llama2-13B-IT Decoder-based Yes 13B
Llama2-70B-IT Decoder-based Yes 70B
Llama3-8B Decoder-based No 8B $\left.\begin{array}[]{@{}r@{}}\text{15T+ tokens}\\ \text{15T+ tokens}\\ \text{15T+ tokens}\end{array}\right\}$ Publicly available online data (details unknown, code is 4x larger than Llama2)
Llama3-8B-IT Decoder-based No 8B
Llama3-70B Decoder-based No 70B
Phi3-mini Decoder-based Yes 3.8B $\left.\begin{array}[]{@{}r@{}}\text{4.9T tokens}\\ \text{4.9T tokens}\\ \text{4.9T tokens}\end{array}\right\}$ High-quality materials including educational data, textbook-like generated text, high-quality chats
Phi3-small Decoder-based Yes 7B
Phi3-medium Decoder-based Yes 14B
†Specify whether the model is an instruction-tuned model or not
*ALBERT shares parameter between token embeddings and transformer layers to compress parameters.
@@ F.3 Evaluation results on all PLMs based on BELIEFs
We present all evaluation results and their computational costs in this section. In Table 17, we report the full-scale experiments using all the prompts provided by MyriadLAMA. This includes PLMs with 8B parameters or fewer, such as BERTbase, BERTlarge, BERTwwm, ALBERTbase, ALBERTlarge, Llama2-7B, Llama3-8B, Llama2-7B-IT, Llama3-8B-IT, Phi3-mini (3.8B), and Phi3-small (7B). For decoder-based models, we conduct experiments on 4 types of ICL settings.
@@
Furthermore, we display the calibration level between accuracy and confidence for a straightforward inspection on Ovconf metrics. We show the calibration figures of models with full-scale experiments in Figure 7 and experiments with partial prompts in Figure 8.
@@
4-random .4816 .2250 .0270 .2312 -.0894 .8247 47.8h
4-relation .6286 .1221 .0150 .3753 -.1335 .9060 47.8h
4-template .6616 .0294 .0036 .4163 -.0933 .9299 47.8h
Llama2-7B-IT zero-shot .2925 .1980 .0253 .1151 .2605 .9069 46.4h
4-random .4334 .1958 .0229 .2128 .2410 .9081 47.8h
4-relation .5576 .0791 .0092 .3341 .1900 .9314 47.8h
4-template .5896 .0439 .0050 .3687 .2061 .9380 47.8h
@@
4-random .5205 .2033 .0273 .2156 -.0796 .8147 82.1h
4-relation .6871 .1236 .0156 .3659 -.0783 .9071 82.1h
4-template .7268 .0220 .0026 .4015 -.0582 .9187 82.1h
Llama3-8B-IT zero-shot .3578 .2213 .0262 .1660 .1402 .7925 44.9h
4-random .4290 .2068 .0222 .2137 .1038 .8511 82.1h
4-relation .5727 .0731 .0092 .3239 .0760 .9140 82.1h
4-template .6508 .0372 .0040 .3727 .0800 .9331 82.1h
@@
(d) Llama2-7B-IT with 4 ICL settings
(e) Llama3-8B with 4 ICL settings
(f) Llama3-8B-IT with 4 ICL settings
(g) Phi3-mini with 4 ICL settings
(h) Phi3-small with 4 ICL settings
@@
(b) Llama2-7B-IT with 4 ICL settings
(c) Llama2-13B with 2 ICL settings
(d) Llama2-13B-IT with 2 ICL settings
(e) Llama2-70B with 2 ICL settings
(f) Llama2-70B-it with 2 ICL settings
(g) Llama3-8B with 2 ICL settings
(h) Llama3-8B-IT with 2 ICL settings
(i) Llama3-70B with 2 ICL settings
(j) Phi3-mini with 4 ICL settings
(k) Phi3-small with 4 ICL settings
(l) Phi3-medium with 2 ICL settings